OpenStudy (anonymous):

check all that apply: If sin theta = 15/17 then: a. tan theta = 15/8 b. csc theta = 17/15 c. coc theta = 17/8 d. sec theta = 17/8

OpenStudy (anonymous):

so just b?

OpenStudy (wolf1728):

If sin theta = 15/17 then opp=15 and hyp =17 and adj = 17² - 15² = 8² adjacent = 8 then tan theta = 15/8 csc theta = hyp/opp = 17/15 coc ??? theta = ??? sec theta = hyp/adj = 17/8 So, they all apply except I don't know what coc is. You don't need to know this but looking up the arc sin of theta gives us the angle 61.928 degrees.

OpenStudy (anonymous):

I think he meant cot

OpenStudy (anonymous):

so cos theta wouldn't apply?

OpenStudy (wolf1728):

Then cot theta = adj / opp = 8/15 so cos theta does NOT apply.

OpenStudy (anonymous):

Nope it doesn't :)

OpenStudy (wolf1728):

or does coc mean cotangent or cotangent?

OpenStudy (anonymous):

I have no clue what coc means lol

OpenStudy (anonymous):

can you help me out with one more please?

OpenStudy (wolf1728):

well okay - but what about coc = 17/8

OpenStudy (anonymous):

oh sorry that was a typo :p

OpenStudy (jdoe0001):

@josereyes so.. what did you get for "a", "b" and "c"?

OpenStudy (wolf1728):

a b and d apply

OpenStudy (anonymous):

a survey measures the angle of elevation to a point on a mountain that is 5 miles away to be 12 degrees. the vertical change in elevation from the point where the surveyor is standing to the point on the mountain is (how many) miles.

OpenStudy (anonymous):

I'm having a hard time with this one.

OpenStudy (wolf1728):

Okay I'm working on a diagram

OpenStudy (anonymous):

Ok

OpenStudy (wolf1728):

So, I guess they are asking for the height of the mountain (the vertical change in elevation)?

OpenStudy (anonymous):

Yes i believe so.

OpenStudy (wolf1728):

Okay well can you figure out what to do about figuring it out? The mountain would be the opposite side and the 5 miles is the adjacent side.

OpenStudy (anonymous):

so would i do tan(12) = x/5

OpenStudy (wolf1728):

Yes but remember you want to solve for x so x = 5 * tan(12°)

OpenStudy (anonymous):

i got 1.06 but that doesn't make sense.

OpenStudy (wolf1728):

I get 0.21256 for tan 12° and that times 5 = 1.0628 That makes sense. And what is the 1.0628 (what units)?

OpenStudy (anonymous):

miles

OpenStudy (wolf1728):

Very good and if the answer was to be in feet you would multiply that by what?

OpenStudy (anonymous):

wait, so the answer is 1.06 miles? and i don't know what would you multiply it by?

OpenStudy (wolf1728):

there are 5,280 feet in a mile so?

OpenStudy (anonymous):

in feet it would be 5596.8

OpenStudy (wolf1728):

or 5,611.584 if you use all the decimals in 1.0628

OpenStudy (anonymous):

Oh alright thanks for your help

OpenStudy (wolf1728):

u r welcome :-)